PROBLEM 3. Find the radius of a cone of revolution of slant height 6 inches in which the total area is twice the lateral area.
FINAL ANSWER: _______________________________________
1
Expert's answer
2020-12-15T02:18:13-0500

Slant Height is l= 6 inches. Let "r" be the radius of cone.

according the question, two times of lateral area of cone "S=\pi r l" is equal to total area "A= \pi r l + \pi r^2" .

"2S =A\\
\Rightarrow2\pi r l=\pi r l +\pi r^2\\
\Rightarrow l=r"

hence radius is equal to slant height which is 6 inches.